Calculate the threshold wavelength, in nm, of a metal if that metal emits an electron with a kinetic energy of 6.32 x 10-19 J when illuminated with light that has a frequency of 1.97 x 1015 Hz. 


1
Expert's answer
2021-04-22T07:28:31-0400

"KE=\dfrac{hc}{\lambda}-W_o"

"W_o=h\nu-KE"

"W_o=(6.63\times10^{-34}\times1.97\times10^{15})-(6.32\times10^{-19})"

"W_o=6.74\times10^{-19}\space J"


For threshold wavelength"(\lambda)" ,

"\dfrac{hc}{\lambda}=W_o"

"\lambda=\dfrac{hc}{W_o}=\dfrac{6.63\times10^{-34}\times3\times10^8}{6.74\times10^{-19}}"

"\lambda=2.95\times10^{-7}\space m"

"\lambda=295\space nm"